# Role of interaction anisotropy in polymer cononsolvency: insights from the Flory-Huggins-Potts framework

### v2

@@ 3.2.2 Chain collapse by preferential mixing of solvent and cosolvent
Chain collapse from an extended solvated polymer to a dry, solvent-excluded globule is observed in systems where solvent-cosolvent interactions are strongly favorable ($\mathcal{R}_{\text{sc}}$). In the isotropic case ($\mathcal{R}_{\text{sc}}^{\circ}$, blue circles), the collapse coincides with nearly equimolar mixtures at which entropy of mixing is maximized, resulting in a symmetric $R_{g}$ profile as a function of cosolvent mole fraction (Fig. 3a); the collapse is likewise evident by the enrichment in monomer-monomer interactions (Fig. 3b), which is also nearly symmetric. The resulting globule is dry, as indicated by the depletion in excess monomer-cosolvent interactions(Fig. 3c). Furthermore, the chain is maximally collapsed at $x_{\text{c}}\approx 0.5$, which aligns with the excess number of solvent-cosolvent interactions (Fig. 3d). This dry globule and congruence of polymer collapse with enrichment in solvent-cosolvent interactions are defining characteristics of cononsolvency driven by solvent-cosolvent interactions.
@@
However, there are notable differences in the anisotropic case ($\mathcal{R}_{\text{sc}}^{\curlywedge}$, orange triangles). In particular, the maximal chain collapse shifts to lesser cosolvent fractions–in this case, around $x_{\text{c}}\approx 0.2$ (Fig. 3a). This is consistent with the maximum in the number of monomer-monomer contacts (Fig. 3b), but it is interestingly inconsistent with maximum enrichemnt of solvent-cosolvent interactions, which remains at equimolar concentrations. The collapsed state of the polymer remains dry, largely excluding both solvent and cosolvent. Collectively, these observations highlight that the distinguishing feature between these two systems is the composition at which the polymer is maximally collapsed. Whereas both systems should effectively share the same macroscopic phase behavior (at the given temperature), systems with strong anisotropy will be characterized by a more asymmetric $R_{g}$ profile as a function of cosolvent mole fraction.
@@
Fig.  4: Analysis and comparison of the solvation environment of systems $\mathcal{R}_{\text{sc}}^{\circ}$ and $\mathcal{R}_{\text{sc}}^{\curlywedge}$. (a) Fraction of bridging cosolvents $\Phi_{\text{b}}$, (b) fraction of mediating cosolvents $\Phi_{\text{m}}$, (c) energetic contribution per cosolvent particle in the solvation shell $\bar{E}_{b}$, and (d) the energetic contribution per cosolvent particle in the solvation shell $\bar{E}_{m}$ as cosolvent fraction $x_{c}$ is varied from 0 to 111. Horizontal axis labels are share betwen panels (a) and (c) as well as (b) and (d). The white arrows highlight a marked difference in the magnitude of intensive energies. Error bars correspond to the standard error of the mean and are generally smaller than the symbol size.
@@
To better understand what drives this asymmetric collapse, we perform a configurational and energetic analysis of the polymer solvation shell. In both systems, on trace addition of cosolvent, nearly all cosolvent particles in the solvation shell act as mediating cosolvents, coordinating interactions between monomer and solvent particles, with the remainder acting as bridging cosolvents (Fig. 4a,b). The relative populations of cosolvent particles and how they interact with the polymer is thus not particularly distinctive. However, the relative energetic contributions from these groups does distinguish the two systems effectively. Each bridging and mediating cosolvent in the anisotropic system provides substantially stronger stabilizing interactions to the polymer over certain composition ranges (Fig. 4c,d). The difference in stabilization between isotropic and anisotropic systems is largest around $x_{c}\in[0.2,0.4]$, thereby providing a substantial enthalpic driving force at low cosolvent fractions.
@@
Fig.  5: A comparison of conformational characteristics and environment of a polymer from $\mathcal{R}_{\text{pc}}^{\circ}$ (blue circles) and $\mathcal{R}_{\text{pc}}^{\curlywedge}$ (orange triangles). (a) Normalized single-chain $R_{\text{g}}$ and normalized excess (b) monomer-monomer, (c) solvent-cosolvent, and (d) monomer-cosolvent interactions as cosolvent fraction $x_{c}$ is varied from 0 to 111. In (a), the dashed horizontal line is a guide for the reference to a maximally compact polymer. Horizontal axis labels are share betwen panels (a) and (c) as well as (b) and (d). In (c) and (d), the white arrow shows a marked shift in the trendlines. Error bars correspond to the standard error of the mean and are generally smaller than the symbol size.
@@ 3.2.3 Chain collapse by preferential adsorption of cosolvent
In a regime dominated by polymer–cosolvent affinity ($\mathcal{R}_{\mathrm{pc}}$), most signatures of microscopic cononsolvency are effectively equivalent between systems with equivalent $\chi^{\text{FHP}}_{ij}$, irrespective of whether such parameters arise from purely isotropic or include anisotropic terms. For example, both isotropic ($\mathcal{R}_{\mathrm{pc}}^{\circ}$, blue circles) and anisotropic ($\mathcal{R}_{\mathrm{pc}}^{\curlywedge}$ orange triangles) systems display asymmetric cosolvent-induced coil–globule transitions, as monitored by $R_{\text{g}}$ (Fig. 5a). Notably, this asymmetric profile may be reminiscent of that found for the anisotropic system system in Fig. 3a. Nevertheless, by comparing the excess number of monomer-monomer interactions in Fig. 3b versus those in Fig. 5b, one can infer that the collapsed chain induced by polymer-cosolvent interactions is relatively wet. In particular, the globules are found to be cosolvent-laden globule, as evidenced by an excess of cosolvent–monomer interactions (Fig. 5c). Between isotropic and anisotropic systems, this enrichment is slightly stronger in the latter. The simulations with anisotropic interactions display more significant solvent-cosolvent interactions at intermediate cosolvent fractions (Fig. 5d). However, this enhanced mixing is of little consequence to the polymer conformational behavior and simply arises to due $\Delta_{\text{sc}}$ being negative. Because of the strong enrichment in polymer-cosolvent interactions, we mechanistically refer to behavior in this regime as preferential adsorption.
@@
Thus, cononsolvency via preferential adsorption exhibits qualitatively similar behavior for both isotropic and anisotropic interactions. In both cases, cosolvent particles percolate the gyration volume, expel solvent, and intercalate between monomer segments to form wet globules. When interactions are only isotropic, this produces abrupt collapse where cosolvent mediates monomer-solvent interactions; the addition of more cosolvent leads to fewer such interactions, causing gradual re-expansion. Anisotropic interactions intensify these physics through enhanced cosolvent intercalation and stronger mediating interactions that persist throughout the transition. While the underlying physical drives remain effectively identical, structural and compositional correlations in globules induced by polymer-cosolvent affinity may differ for sufficiently strong anisotropic interactions.
Fig.  6: Analysis and comparison of the solvation environment of systems $\mathcal{R}_{\text{pc}}^{\circ}$ and $\mathcal{R}_{\text{pc}}^{\curlywedge}$. (a) Fraction of bridging cosolvents $\Phi_{\text{b}}$, (b) fraction of mediating cosolvents $\Phi_{\text{m}}$, (c) energetic contribution per cosolvent particle in the solvation shell $\bar{E}_{b}$, and (d) the energetic contribution per cosolvent particle in the solvation shell $\bar{E}_{m}$ as cosolvent fraction $x_{c}$ is varied from 0 to 111. Horizontal axis labels are share betwen panels (a) and (c) as well as (b) and (d). In (c) and (d), the white arrow shows a marked shift in the intensive energy. Error bars correspond to the standard error of the mean and are generally smaller than the symbol size.
@@ 4 Conclusions
We investigated how orientational interactions influence cononsolvency phenomena in ternary polymer solutions. Using the Flory-Huggins-Potts (FHP) framework enabled systematic comparison of systems with identical effective $\chi$ parameters but different underlying interaction types, including those with orientation-dependent energetic contributions. After identifying phase-separation regimes driven by either solvent-cosolvent or polymer-cosolvent affinity, we contrasted systems with the same effective $\chi$ parameters but achieved either purely through isotropic interactions or with inclusion of anisotropic interactions, specifically with cosolvent species. This approach enabled controlled comparisons between systems with identical mean-field drivers, isolating how anisotropic interactions affect microscopic physics despite equivalent macroscopic phase behavior. Analysis focused on cosolvent-induced coil-globule transitions. This work complements prior extensive literature on consolvency based on Flory-Huggins (FH) theory.
@@
Our analysis revealed that interaction anisotropy has variable importance depending on the dominant interaction type driving cononsolvency. For systems driven by strong solvent-cosolvent interactions, orientation-dependent interactions qualitatively altered coil-globule transitions, inducing collapse at lower cosolvent fractions than isotropic systems with equivalent mean-field parameters. This enhanced, asymmetric collapse was attributed to stronger cosolvent-mediated interactions, where cosolvent acts as a surfactant favorably interacting with both polymer and solvent.44 This asymmetric collapse notably resembled that observed in polymer-cosolvent affinity-driven systems. However, whereas solvent-cosolvent affinity produced dry, solvent-depleted globules, those produce by polymer-cosolvent affinity were more wet with embedded cosolvent; this distinction offers a pathway to distinguish these regimes at the microscopic level. When comparing isotropic and anisotropic systems both dominated by polymer-cosolvent affinity, anisotropic interactions primarily increase bridging-type cosolvent configurations and potentially enhance globule stability, without fundamentally altering the mechanism.
@@
This work highlights several implications and opportunities for future inquiry. Discriminating among cononsolvency mechanisms proves challenging when examining only macroscopic phase behavior or microscopic single-chain conformations in isolation. However, complementary analysis at both scales suggests that preferential mixing, anisotropically-influenced preferential mixing, and preferential adsorption mechanisms can be distinguished. While investigating the role of anisotropy interactions via experiment remains nontrivial, our results suggest that temperature-dependent characterization across compositions may be informative. While the FHP framework provides physically grounded parameters enabling precise experimental fitting,66 this has so far relied on macroscopic phase behavior. Connecting FHP parameters to molecular simulations or microscopic experimental observables would clarify when orientation-dependent interactions emerge in real systems and their representation within FHP. Meanwhile, we propose that studies employing systematic molecular modifications—such as adding electron-withdrawing substituents to cosolvents to tune the strength of hydrogen-bonding—may offer tangible routes to probe anisotropic effects without requiring FHP abstraction.
@@ Acknowledgements
  • Wang et al. 2019 X. Wang, H. Huang, H. Liu, F. Rehfeldt, X. Wang and K. Zhang, Macromolecular Chemistry and Physics, 2019, 220, xx–yy.
@@
  • Canchi and Garc´ıa 2013 D. R. Canchi and A. E. García, Annual Review of Physical Chemistry, 2013, 64, 273–293.
@@
  • Rodr´ıguez-Cabello et al. 2016 J. C. Rodríguez-Cabello, F. J. Arias, M. A. Rodrigo and A. Girotti, Advanced Drug Delivery Reviews, 2016, 97, 85–100.
@@
  • Pérez-Ram´ırez et al. 2019 H. A. Pérez-Ramírez, C. Haro-Pérez, E. Vázquez-Contreras, J. Klapp, G. Bautista-Carbajal and G. Odriozola, Physical Chemistry Chemical Physics, 2019, 21, 5106–5116.
@@
  • Bharadwaj et al. 2020 S. Bharadwaj, D. Nayar, C. Dalgicdir and N. F. A. van der Vegt, Communications Chemistry, 2020, 3, xx–yy.
  • Dudowicz et al. 2015 J. Dudowicz, K. F. Freed and J. F. Douglas, The Journal of Chemical Physics, 2015, 142, xx–yy.
@@
  • Bharadwaj et al. 2021 S. Bharadwaj, D. Nayar, C. Dalgicdir and N. F. A. van der Vegt, The Journal of Chemical Physics, 2021, 154, xx–yy.
  • Mukherji et al. 2014 D. Mukherji, C. M. Marques and K. Kremer, Nature Communications, 2014, 5, xx–yy.
@@
  • van der Vegt and Rodr´ıguez-Ropero 2017 N. F. A. van der Vegt and F. Rodríguez-Ropero, Soft Matter, 2017, 13, 2289–2291.
@@
  • Swathi et al. 2022 P. V. Swathi, U. Abdulkareem, T. R. Kartha and V. Madhurima, ChemistrySelect, 2022, 7, xx–yy.

### v3

@@ 3.2.2 Chain collapse by preferential mixing of solvent and cosolvent
Chain collapse from an extended solvated polymer to a dry, solvent-excluded globule is observed in systems where solvent-cosolvent interactions are strongly favorable ($R_{\text{sc}}$). In the isotropic case ($R_{\text{sc}}^{\circ}$, blue circles), the collapse coincides with nearly equimolar mixtures at which entropy of mixing is maximized, resulting in a symmetric $R_{g}$ profile as a function of cosolvent mole fraction (Fig. 3a); the collapse is likewise evident by the enrichment in monomer-monomer interactions (Fig. 3b), which is also nearly symmetric. The resulting globule is dry, as indicated by the depletion in excess monomer-cosolvent interactions (Fig. 3c). The excess number of monomer-cosolvent contacts remain small, indicating behavior close to that expected from ideal mixing and thus limited preferential interaction between cosolvent and polymer. Furthermore, the chain is maximally collapsed at $x_{\text{c}}\approx 0.5$, which aligns with the excess number of solvent-cosolvent interactions (Fig. 3d). This dry globule and congruence of polymer collapse with enrichment in solvent-cosolvent interactions are defining characteristics of cononsolvency driven by solvent-cosolvent interactions.
@@
However, there are notable differences in the anisotropic case ($R_{\text{sc}}^{\curlywedge}$, orange triangles). In particular, the maximal chain collapse shifts to lesser cosolvent fractions–in this case, around $x_{\text{c}}\approx 0.3$ (Fig. 3a). This is consistent with the maximum in the number of monomer-monomer contacts (Fig. 3b), but it is interestingly inconsistent with maximum enrichemnt of solvent-cosolvent interactions, which remains at equimolar concentrations. The collapsed state of the polymer remains dry, largely excluding both solvent and cosolvent. Collectively, these observations highlight that the distinguishing feature between these two systems is the composition at which the polymer is maximally collapsed. Whereas both systems should effectively share the same macroscopic phase behavior (at the given temperature), systems with strong anisotropy will be characterized by a more asymmetric $R_{g}$ profile as a function of cosolvent mole fraction.
@@
Fig. 4: Analysis and comparison of the solvation environment of systems $R_{\text{sc}}^{\circ}$ and $R_{\text{sc}}^{\curlywedge}$. (a) Fraction of bridging cosolvents $\Phi_{\text{b}}$, (b) fraction of mediating cosolvents $\Phi_{\text{m}}$, (c) energetic contribution per cosolvent particle in the solvation shell $\bar{E}_{b}$, and (d) the energetic contribution per cosolvent particle in the solvation shell $\bar{E}_{m}$ as cosolvent fraction $x_{c}$ is varied from 0 to 11. Horizontal axis labels are share betwen panels (a) and (c) as well as (b) and (d). The white arrows highlight a marked difference in the magnitude of intensive energies. Error bars correspond to the standard error of the mean and are generally smaller than the symbol size.
@@
To better understand what drives this asymmetric collapse, we perform a configurational and energetic analysis of the polymer solvation shell. In both systems, on trace addition of cosolvent, nearly all cosolvent particles in the solvation shell act as mediating cosolvents, coordinating interactions between monomer and solvent particles, with the remainder acting as bridging cosolvents (Fig. 4a,b). The relative populations of cosolvent particles and how they interact with the polymer is thus not particularly distinctive. However, the relative energetic contributions from these groups does distinguish the two systems effectively. Each bridging and mediating cosolvent in the anisotropic system provides substantially stronger stabilizing interactions to the polymer over certain composition ranges (Fig. 4c,d). The difference in stabilization between isotropic and anisotropic systems is largest around $x_{c}\in[0.2,0.4]$, thereby providing a substantial enthalpic driving force at lower cosolvent fractions and shifting the collapse from $x_{\text{c}}=0.5$ in the isotropic case to $x_{\text{c}}=0.3$ in the anisotropic case.
@@
Fig. 5: A comparison of conformational characteristics and environment of a polymer from $R_{\text{pc}}^{\circ}$ (blue circles) and $R_{\text{pc}}^{\curlywedge}$ (orange triangles). (a) Normalized single-chain $R_{\text{g}}$ and normalized excess (b) monomer-monomer, (c) solvent-cosolvent, and (d) monomer-cosolvent interactions as cosolvent fraction $x_{c}$ is varied from 0 to 11. In (a), the dashed horizontal line is a guide for the reference to a maximally compact polymer. Horizontal axis labels are share betwen panels (a) and (c) as well as (b) and (d). In (c) and (d), the white arrow shows a marked shift in the trendlines. Error bars correspond to the standard error of the mean and are generally smaller than the symbol size.
@@ 3.2.3 Chain collapse by preferential adsorption of cosolvent
In a regime dominated by polymer–cosolvent affinity ($R_{\mathrm{pc}}$), most signatures of microscopic cononsolvency are effectively equivalent between systems with equivalent $\chi^{\text{FHP}}_{ij}$, irrespective of whether such parameters arise from purely isotropic or include anisotropic terms. For example, both isotropic ($R_{\mathrm{pc}}^{\circ}$, blue circles) and anisotropic ($R_{\mathrm{pc}}^{\curlywedge}$ orange triangles) systems display asymmetric cosolvent-induced coil–globule transitions, as monitored by $R_{\text{g}}$ (Fig. 5a). Notably, this asymmetric profile may be reminiscent of that found for the anisotropic system system in Fig. 3a. Nevertheless, by comparing the excess number of monomer-monomer interactions in Fig. 3b versus those in Fig. 5b, one can infer that the collapsed chain induced by polymer-cosolvent interactions is relatively wet. In particular, the globules are found to be cosolvent-laden globule, as evidenced by an excess of cosolvent–monomer interactions (Fig. 5c). Between isotropic and anisotropic systems, this enrichment is slightly stronger in the latter. The simulations with anisotropic interactions display more significant solvent-cosolvent interactions at intermediate cosolvent fractions (Fig. 5d). However, this enhanced mixing is of little consequence to the polymer conformational behavior and simply arises to due $\Delta_{\text{sc}}$ being negative. Because of the strong enrichment in polymer-cosolvent interactions, we mechanistically refer to behavior in this regime as preferential adsorption.
@@
Thus, cononsolvency via preferential adsorption exhibits qualitatively similar behavior for both isotropic and anisotropic interactions. In both cases, cosolvent particles percolate the gyration volume, expel solvent, and intercalate between monomer segments to form wet globules. This supports prior observations56 that a polymer chain can collapse without reduction in overall solvent quality. In this case, this arises because the cosolvent intercalates with strong affinity to the polymer chain. When interactions are only isotropic, this produces abrupt collapse where cosolvent mediates monomer-solvent interactions; the addition of more cosolvent leads to fewer such interactions, causing gradual re-expansion. Anisotropic interactions intensify these physics through enhanced cosolvent intercalation and stronger mediating interactions that persist throughout the transition. While the underlying physical drives remain effectively identical, structural and compositional correlations in globules induced by polymer-cosolvent affinity may differ for sufficiently strong anisotropic interactions.
Fig. 6: Analysis and comparison of the solvation environment of systems $R_{\text{pc}}^{\circ}$ and $R_{\text{pc}}^{\curlywedge}$. (a) Fraction of bridging cosolvents $\Phi_{\text{b}}$, (b) fraction of mediating cosolvents $\Phi_{\text{m}}$, (c) energetic contribution per cosolvent particle in the solvation shell $\bar{E}_{b}$, and (d) the energetic contribution per cosolvent particle in the solvation shell $\bar{E}_{m}$ as cosolvent fraction $x_{c}$ is varied from 0 to 11. Horizontal axis labels are share betwen panels (a) and (c) as well as (b) and (d). In (c) and (d), the white arrow shows a marked shift in the intensive energy. Error bars correspond to the standard error of the mean and are generally smaller than the symbol size.
@@ 4 Conclusions
We investigated how orientational interactions influence cononsolvency phenomena in ternary polymer solutions. Using the Flory-Huggins-Potts (FHP) framework enabled systematic comparison of systems with identical effective $\chi$ parameters but different underlying interaction types, including those with orientation-dependent energetic contributions. After identifying phase-separation regimes driven by either solvent-cosolvent or polymer-cosolvent affinity, we contrasted systems with the same effective $\chi$ parameters but achieved either purely through isotropic interactions or with inclusion of anisotropic interactions, specifically with cosolvent species. This approach enabled controlled comparisons between systems with identical mean-field drivers, isolating how anisotropic interactions affect microscopic physics despite equivalent macroscopic phase behavior. Analysis focused on cosolvent-induced coil-globule transitions. This work complements prior extensive literature on cononsolvency based on FH theory.
@@
Our analysis revealed that orientation-dependent interactions have variable importance depending on the dominant interaction type driving cononsolvency. For systems driven by strong solvent-cosolvent interactions, orientation-dependent interactions qualitatively altered coil-globule transitions, inducing collapse at lower cosolvent fractions than isotropic systems with equivalent mean-field parameters. This enhanced, asymmetric collapse was attributed to stronger cosolvent-mediated interactions, where cosolvent acts as a surfactant favorably interacting with both polymer and solvent.47 This asymmetric collapse notably resembled that observed in polymer-cosolvent affinity-driven systems. However, whereas solvent-cosolvent affinity produced dry, solvent-depleted globules, those produce by polymer-cosolvent affinity were more wet with embedded cosolvent; this distinction offers a pathway to distinguish these regimes at the microscopic level. When comparing isotropic and anisotropic systems both dominated by polymer-cosolvent affinity, anisotropic interactions primarily increase bridging-type cosolvent configurations and potentially enhance globule stability, without fundamentally altering the mechanism.
@@
This work highlights several implications and opportunities for future inquiry. Discriminating among cononsolvency mechanisms proves challenging when examining only macroscopic phase behavior or microscopic single-chain conformations in isolation. However, complementary analysis at both scales suggests that preferential mixing, anisotropically-influenced preferential mixing, and preferential adsorption mechanisms can be distinguished. While investigating the role of anisotropy interactions via experiment remains nontrivial, our results suggest that temperature-dependent characterization across compositions may be informative. While the FHP framework provides physically grounded parameters enabling precise experimental fitting,70 this has so far relied on macroscopic phase behavior. Connecting FHP parameters to molecular simulations or microscopic experimental observables would clarify when orientation-dependent interactions emerge in real systems. Meanwhile, we propose that studies employing systematic molecular modifications—such as adding electron-withdrawing substituents or stackable molecular moieties to cosolvents to tune hydrogen-bonding and molecular packing—may offer tangible routes to probe anisotropic effects without requiring FHP abstraction.
@@ Acknowledgements
  • Wang et al. 2019 X. Wang, H. Huang, H. Liu, F. Rehfeldt, X. Wang and K. Zhang, Macromolecular Chemistry and Physics, 2019, 220, .
@@
  • Canchi and García 2013 D. R. Canchi and A. E. García, Annual Review of Physical Chemistry, 2013, 64, 273–293.
@@
  • Rodríguez-Cabello et al. 2016 J. C. Rodríguez-Cabello, F. J. Arias, M. A. Rodrigo and A. Girotti, Advanced Drug Delivery Reviews, 2016, 97, 85–100.
  • Dudowicz et al. 2015 J. Dudowicz, K. F. Freed and J. F. Douglas, The Journal of Chemical Physics, 2015, 142, 214906.
  • Schild et al. 1991 H. G. Schild, M. Muthukumar and D. A. Tirrell, Macromolecules, 1991, 24, 948–952.
@@
  • Pérez-Ramírez et al. 2019 H. A. Pérez-Ramírez, C. Haro-Pérez, E. Vázquez-Contreras, J. Klapp, G. Bautista-Carbajal and G. Odriozola, Physical Chemistry Chemical Physics, 2019, 21, 5106–5116.
@@
  • Bharadwaj et al. 2020 S. Bharadwaj, D. Nayar, C. Dalgicdir and N. F. A. van der Vegt, Communications Chemistry, 2020, 3, 1–7.
@@
  • Oliver et al. 2025 W. W. Oliver, W. M. Jacobs and M. A. Webb, The Journal of Physical Chemistry B, 2025, 129, 9551–9565.
  • Liu et al. 2024 L. Liu, C. Duan and R. Wang, Macromolecules, 2024, 57, 10694–10703.
  • Bharadwaj et al. 2021 S. Bharadwaj, D. Nayar, C. Dalgicdir and N. F. A. van der Vegt, The Journal of Chemical Physics, 2021, 154, 134903.
  • Mukherji et al. 2014 D. Mukherji, C. M. Marques and K. Kremer, Nature Communications, 2014, 5, 1–6.
@@
  • van der Vegt and Rodríguez-Ropero 2017 N. F. A. van der Vegt and F. Rodríguez-Ropero, Soft Matter, 2017, 13, 2289–2291.
@@
  • Swathi et al. 2022 P. V. Swathi, U. Abdulkareem, T. R. Kartha and V. Madhurima, ChemistrySelect, 2022, 7, e202200413.